OpenStudy (anonymous):

can anyone explain a Algebra problem to me? (3Sqrt7-3sqrt3)^2

OpenStudy (anonymous):

I just need to know how the steps to solve it

OpenStudy (anonymous):

thanks Night watcher

OpenStudy (anonymous):

\[(3\sqrt{7}-3\sqrt{3})^{2}\]

OpenStudy (anonymous):

thanks you so much I just need to know how to do the problem like the formula or something

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

Simply this one (3√7) Then simplify this one (3√3) Subtract both answers Then square it

OpenStudy (anonymous):

3sqrt7 and 3sqrt3 squared is how to do it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what do i subtract

OpenStudy (anonymous):

You can find this symbol √ on a calculator. so first type 3 times 7 then hit this button √. After you did that you simplified the first one.

OpenStudy (anonymous):

Then do the same for the second one

OpenStudy (amistre64):

\[(a+b)^2=(a+b)(a+b)=a^2+b^2+2ab\]

OpenStudy (anonymous):

i am stil lost hold on

OpenStudy (anonymous):

Then you'll subtract both of them answers and square them.

OpenStudy (anonymous):

hold on I am lost so srry. Let me catch up

OpenStudy (anonymous):

so you take the (3sqrt7-3sqrt3)^2 converts to 3sqrt7 and 3sqrt3 then I have to subtract what then square what

OpenStudy (anonymous):

@amistre64 can u help

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (amistre64):

i did help .... use the formula.

OpenStudy (anonymous):

ok I am trying. so you take the (3sqrt7-3sqrt3)^2 converts to 3sqrt7 and 3sqrt3 then I have to subtract what then square what is this part right @amistre64

OpenStudy (amistre64):

i would not do that ...

OpenStudy (amistre64):

i would do:\[(a+b)=a^2+b^2+2ab\]

OpenStudy (anonymous):

why?

OpenStudy (amistre64):

sqrt7 and sqrt3 are irrational numbers that have no exact decimal forms to play with.

OpenStudy (anonymous):

ok so that would be umm 3sqrt7+3sqrt3 is that part corrrect

OpenStudy (anonymous):

sorry I am horrible at Algebra

OpenStudy (anonymous):

I am trying hard but my math teacher explains it way to quick and complexx

OpenStudy (amistre64):

\[(~\underbrace{3\sqrt7}_a~+~\underbrace{(-3\sqrt3)}_b~)^2\]

OpenStudy (anonymous):

omg i have that part on paper but i thought it was wrong hold on

OpenStudy (amistre64):

\[(a+b)=a^2+b^2+2ab\] \[(3\sqrt7-3\sqrt3)=(3\sqrt7)^2+(-3\sqrt3)^2+2(3\sqrt7)(-3\sqrt3)\]

OpenStudy (anonymous):

thanks so much i got it Thanks so very very much

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thanks I fanned u and medaled u